# Improving entanglement purification through coherent superposition of roles

### v1

@@ Abstract
Entanglement purification and distillation protocols are essential for harnessing the full potential of quantum communication technologies. Multiple strategies have been proposed to approach and optimize such protocols, most however restricted to Clifford operations. In this paper, we introduce a novel superposed entanglement purification design strategy, leveraging coherent superpositions of the roles of entangled states to enhance purification efficiency. We demonstrate how this approach can be hierarchically integrated with existing entanglement purification strategies, consistently improving protocols performance.
I Introduction
@@
Entanglement purification protocols (EPPs) essentially involve distilling a few high-quality entangled states from a larger number of lower-quality copies, achieved by using only local operations at each involved party. Different strategies have been proposed, including recurrence [17, 18, 19, 20] and hashing-like [17, 19, 21, 22] protocols, that work for bipartite [17, 18, 23, 24, 16, 25, 26] or multipartite [27, 28, 29, 30] entanglement, as well as for qudits [31, 32, 33] or other kinds of entanglement[34, 35, 36, 37, 38]. All of them have in common an inner working mechanism, that consists in transferring information from some copies into other one(s). These copies are subsequently measured to learn some information about the remaining states, eventually increasing their fidelity, either in a probabilistic or a deterministic way.
EPPs have undergone extensive investigation, yielding optimal constructions in some cases [39, 40, 41, 42, 43, 44]. However, concrete protocol proposals are predominantly confined to Clifford operations and therefore fail to explore potential protocol gains beyond Clifford circuits. In this work, we propose an alternative strategy inspired by approaches in diverse contexts [33, 45, 46, 47, 48, 49], enabling the development of enhanced recurrence protocols for entanglement purification. Our method employs a simple yet powerful technique leveraging tools beyond the Clifford group. Fundamentally, the approach involves applying diverse EPP configurations coherently in a superposition, enabling simultaneous execution of these configurations in superposition. This strategy draws motivation from successful implementations in other quantum computation and communication contexts [45, 46, 47, 49, 50, 51]. As we show below, such superpositions can be achieved using a quantum control plane, that allows coherent state role exchanges or complete protocol coherent exchanges. Naturally, this entails increased resource demands, yet we demonstrate how the recursive nature of our approach can consistently outperform existing methods and protocols. In fact, entanglement is required to perform tasks or protocols in coherent superposition, which is simply provided by one (or several) additional copy (copies) of the noisy initial entangled state. Our approach allows one to combine arbitrary existing EPPs, and obtain new protocols operating on slightly more copies, but in many cases with enhanced performance.
@@
This work is structured as follows. In Sec. II, we review the relevant notions and tools utilized in this paper. Sec. III introduces our proposed role-exchanging superposed method, wherein we analyze and compare its performance, including noisy conditions. The extension of these ideas is discussed in Sec. IV. We summarize and conclude in Sec. V.
II Background
@@ II.1 Bell states
II.2 Noisy channels and state fidelity
@@
$\rho=F\left|\Phi_{00}\right\rangle\!\left\langle\Phi_{00}\right|+\frac{1-F}{4}\sum_{i}(\sigma_{i}\otimes\mathbb{1})\left|\Phi_{00}\right\rangle\!\left\langle\Phi_{00}\right|(\sigma_{i}\otimes\mathbb{1}),$ (5)
where $\sigma_{i}=\{\mathbb{1},\sigma_{z},\sigma_{x},\sigma_{y}\}$ for $i=\{0,1,2,3\}$. This equals
@@
with $F=\frac{3q+1}{4}$. This represent a worst-case scenario, since any other state can be brought to this Werner form without altering its fidelity, by means of depolarization techniques [54]. Finally, we also consider amplitude damping (or decay) channel $\xi_{\rm damp}$, given by
@@
II.3 Entanglement purification protocols (EPPs)
@@
Given the conceptual scope of this work, we primarily focus on recurrence EPPs, initially introduced in [17, 55, 18]. In its basic setting, the BBPSSW approach [17] (see Fig. 1a) iteratively operates on two noisy entangled states. Basic two-qubit operations are performed at each site between the copies, effectively transferring some information from one copy to the other. Subsequent measurement of the latter copy enables one to effectively learn some information about the former, resulting in a probabilistic increase in the fidelity of the surviving state. The selection process requires two-way classical communication. The Oxford (or DEJMPS) variant of the protocol [18] includes a local unitary rotation $U_{O}$ given by
$\left|0\right\rangle_{\rm A}\rightarrow\tfrac{1}{\sqrt{2}}(\left|0\right\rangle_{\rm A}-i\left|1\right\rangle_{\rm A}),\quad\left|1\right\rangle_{\rm A}\rightarrow\tfrac{1}{\sqrt{2}}(\left|1\right\rangle_{\rm A}-i\left|0\right\rangle_{\rm A});$ (8)
$\left|0\right\rangle_{\rm B}\rightarrow\tfrac{1}{\sqrt{2}}(\left|0\right\rangle_{\rm B}+i\left|1\right\rangle_{\rm B}),\quad\left|1\right\rangle_{\rm B}\rightarrow\tfrac{1}{\sqrt{2}}(\left|1\right\rangle_{\rm B}+i\left|0\right\rangle_{\rm B}).$ (9)
In addition, the so-called P1-P2 variation of both EPPs includes an adaptive application of the local Hadamard rotation, $H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$. This allows the definition of two subroutines, P1 and P2, which operate with and without the Hadamard gate, respectively, and detect different kinds of errors. These subroutines can be iteratively alternated or optimally selected to further enhance the performance of the protocols.
@@
where obtained copies need to have a fidelity above some pre-defined threshold value $F_{0}$. For recurrence protocols with a single copy as output, the obtained number of copies is given by $\prod_{i}P_{i}$, where $P_{i}$ is the success probability of the $i^{\rm th}$ iteration.
@@
II.4 Superposition of quantum channels
@@
Superposition of different tasks is achieved by introducing a quantum control register in a superposition state, which acts as a control to exchange roles of states or tasks via a controlled-SWAP operation, also known as Fredkin gate [57, 58], described as
@@
III Superposed EPPs: role exchange
@@ III.1 Basic protocol
    $\left|\Phi_{00}\right\rangle\left\langle\Phi_{00}\right|_{\rm A_{\rm c}B_{\rm c}}$ $\otimes\left(\sum_{i,j=0}^{1}{p}_{i,j}\Sigma_{i,j}\left|\Phi_{00}\right\rangle\left\langle\Phi_{00}\right|_{\rm A_{1}B_{1}}\Sigma^{\dagger}_{i,j}\right)$ (13)
@@
    where we can assume without loss of generality that $\Sigma_{ij}\equiv\mathbb{1}_{\rm A}\otimes{\sigma_{ij}}_{\rm B}$, where $\{\sigma_{00},\sigma_{01},\sigma_{10},\sigma_{11}\}=\{\mathbb{1},\sigma_{z},\sigma_{x},\sigma_{y}\}$, such that $\Sigma_{ij}\left|\Phi_{00}\right\rangle=\left|\Phi_{ij}\right\rangle$ maps all the basis Bell states.
@@
    $\frac{1}{2}\left|{ss}\right\rangle\left\langle{ss}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{i,j}\right\rangle\left\langle\Phi_{i,j}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{q,r}\right\rangle\left\langle\Phi_{q,r}\right|_{\rm A_{2}B_{2}}\right)$ (14)
@@
    $+\frac{1}{2}\left|{11}\right\rangle\left\langle{00}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{q,r}\right\rangle\left\langle\Phi_{i,j}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{i,j}\right\rangle\left\langle\Phi_{q,r}\right|_{\rm A_{2}B_{2}}\right).$
@@
    $\left|{ss}\right\rangle\left\langle{ss}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{i\oplus q,j}\right\rangle\left\langle\Phi_{i\oplus q,j}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{q,r\oplus j}\right\rangle\left\langle\Phi_{q,r\oplus j}\right|_{\rm A_{2}B_{2}}\right)+$ (15)
    $\frac{1}{2}\left|{00}\right\rangle\left\langle{11}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{i\oplus q,j}\right\rangle\left\langle\Phi_{q\oplus i,r}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{q,r\oplus j}\right\rangle\left\langle\Phi_{i,j\oplus r}\right|_{\rm A_{2}B_{2}}\right)+$
    $\frac{1}{2}\left|{11}\right\rangle\left\langle{00}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{q\oplus i,r}\right\rangle\left\langle\Phi_{i\oplus q,j}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{i,j\oplus r}\right\rangle\left\langle\Phi_{q,r\oplus j}\right|_{\rm A_{2}B_{2}}\right).$
@@
    Subsequently, another controlled-SWAP (cSWAP) operation can be implemented, although this step is not necessary. Finally, the second copy is measured locally in the $Z$ basis at each site. If parties $A$ and $B$ obtain the same outcome ($j\oplus r=0$), the first state is retained. The only components that survive the measurement are then the ones steming from $\left({p}_{q,j}\left|\Phi_{i,0}\right\rangle\left\langle\Phi_{q,0}\right|_{\rm A_{2}B_{2}}\right)$. Simultaneously, the control state is measured locally in the $X$ basis by $\rm A$ and $\rm B$, post-selecting the desired outcomes as stated in the following.
@@
where $P=\sum_{j}{p}^{2}_{jj}$ is a normalization constant that equals the success probability.
@@
To address this, we turn our attention to a realistic scenario: a $3\rightarrow 1$ protocol involving a control copy identically noisy as the others, as analyzed in Fig. 3. We show how the role-exchanging superposed entanglement purification approach introduced above, can significantly surpass single selection protocols [17, 23, 18], and also more elaborated ones such as the double selection protocol, Fig. 1, [39], which is considered among the most efficient EPPs modifications using $3$ raw EPR copies [41]. In addition, we consider the P1-or-P2 variants of all the protocols (see Figs. 1 - 2), meaning that the best subroutine (P1 or P2) is selected in each iteration. We analyze the performance for different kinds of initial states. The plots show the increase in fidelity with the number of steps, where one observes that the superposed protocol performs better than the elementary ones.
@@
Fig. 3: (a) Fidelity evolution as a function of the number iterations for the BBPSSW P1-or-P2 EPPs with initial states affected (both sides) by amplitude damping noise of strength $q=0.6$. (b) Infidelity evolution as a function of the number iterations for the different P1-or-P2 Oxford variation protocols with initial states of fidelity $F_{0}=0.6$ and dephasing noise (see Appendix Appendix A for a detailed analysis of each protocol for correcting Pauli noises). In all cases, depolarizing the initial states into Werner form decreases the overall protocols performances.
III.2 Imperfect operations

### v2

@@ Abstract
Entanglement purification and distillation protocols are essential for harnessing the full potential of quantum communication technologies. Multiple strategies have been proposed to approach and optimize such protocols, most however restricted to Clifford operations. In this paper, we introduce a superposed entanglement purification design strategy, leveraging coherent superpositions of the roles of entangled states to enhance purification efficiency, defining a new family of non-Clifford distillation protocols. We demonstrate how this approach can be hierarchically integrated with existing entanglement purification strategies, consistently improving protocols performance.
1 Introduction
@@
Entanglement purification protocols (EPPs) essentially involve distilling a few high-quality entangled states from a larger number of lower-quality copies, achieved by using only local operations at each involved party. Different strategies have been proposed, including recurrence [17, 18, 19, 20] and hashing-like [17, 19, 21, 22] protocols, that work for bipartite [17, 18, 23, 24, 16, 25, 26, 27, 28] or multipartite [29, 30, 31, 32] entanglement, as well as for qudits [33, 34, 35] or other kinds of entanglement [36, 37, 38, 39, 40]. All of them have in common an inner working mechanism, that consists in transferring information from some copies into other one(s). These copies are subsequently measured to learn some information about the remaining states, eventually increasing their fidelity, either in a probabilistic or a deterministic way.
EPPs have undergone extensive investigation, yielding optimal constructions in some cases [41, 42, 43]. However, concrete protocol proposals are predominantly confined to Clifford operations and only a few efforts have explored potential advantages of protocols beyond Clifford circuits [44, 45, 46, 47], largely due to the complexity of optimizing and identifying suitable constructions within larger classes of operations.
In this work, we propose an alternative strategy inspired by approaches from diverse quantum information contexts [35, 48, 49, 50, 51, 52], exploring the unique advantages of channel superposition when applied to basic existing entanglement distillation strategies. Fundamentally, this approach enables diverse EPP configurations to operate simultaneously in coherent superposition. As a side effect of our approach, we naturally introduce a new family of non-Clifford entanglement purification protocols. These protocols can simplify the challenge of searching for solutions in expansive operation sets, particularly those that lie outside the Clifford group.
As we demonstrate, superposition can be achieved using a quantum control plane, inherently provided by one or more additional copies of the noisy initial entangled state, ensuring the resource overhead remains practical. This control plane facilitates coherent state-role exchanges or complete protocol exchanges in superposition. While this increases resource demands, we show that the recursive nature of our approach consistently outperforms existing methods and protocols. Importantly, the observed performance enhancements are not only significant under ideal conditions but also robust in noisy operational regimes and for different kinds of initial states.
Our method provides a versatile framework for combining arbitrary existing EPPs to construct distillation protocols that can operate with enhanced performance. Moreover, this strategy effectively introduces a direct way to design alternative families of entanglement purification protocols, which are not limited to specific cases. In particular, one can apply our approach to any known entanglement purification protocol –thereby improving its performance–, but also to ones which will be developed in the future.
@@
This work is structured as follows. In Sec. 2, we review the relevant notions and tools utilized in this paper. Sec. 3 introduces our proposed role-exchanging superposed method, wherein we analyze and compare its performance, including noisy conditions. The extension of these ideas is discussed in Sec. 4. We summarize and conclude in Sec. 5.
2 Background
@@ 2.1 Bell states
2.2 Noisy channels and state fidelity
@@
$\rho=F\left|\Phi_{00}\right\rangle\!\left\langle\Phi_{00}\right|+\frac{1-F}{3}\sum_{i}(\sigma_{i}\otimes\mathbb{1})\left|\Phi_{00}\right\rangle\!\left\langle\Phi_{00}\right|(\sigma_{i}\otimes\mathbb{1}),$ (5)
where $\sigma_{i}=\{\sigma_{z},\sigma_{x},\sigma_{y}\}$ for $i=\{1,2,3\}$. This equals
@@
with $F=\frac{3q+1}{4}$. This represent a worst-case scenario (white noise), since any other state can be brought to this Werner form with local operation and without altering its fidelity (although partially destroying some entanglement), by means of depolarization techniques [55]. Finally, we also consider amplitude damping (or decay) channel $\xi_{\rm damp}$, given by
@@
2.3 Entanglement purification protocols (EPPs)
@@
Given the conceptual scope of this work, we primarily focus on recurrence EPPs, initially introduced in [17, 18]. In its basic setting, the BBPSSW approach [17] (see Fig. 1a) iteratively operates on two noisy entangled states. Basic two-qubit operations are performed at each site between the copies, effectively transferring some information from one copy to the other. Subsequent measurement of the latter copy enables one to effectively learn some information about the former, resulting in a probabilistic increase in the fidelity of the surviving state. The selection process requires two-way classical communication. The Oxford (or DEJMPS) variant of the protocol [18] includes a local unitary rotation $U_{O}$ given by
$\left|0\right\rangle_{\rm A}\rightarrow\tfrac{1}{\sqrt{2}}(\left|0\right\rangle_{\rm A}-i\left|1\right\rangle_{\rm A}),\left|1\right\rangle_{\rm A}\rightarrow\tfrac{1}{\sqrt{2}}(\left|1\right\rangle_{\rm A}-i\left|0\right\rangle_{\rm A}),$ (8)
$\left|0\right\rangle_{\rm B}\rightarrow\tfrac{1}{\sqrt{2}}(\left|0\right\rangle_{\rm B}+i\left|1\right\rangle_{\rm B}),\left|1\right\rangle_{\rm B}\rightarrow\tfrac{1}{\sqrt{2}}(\left|1\right\rangle_{\rm B}+i\left|0\right\rangle_{\rm B}).$ (9)
In addition, the so-called P1-P2 variation of both EPPs includes an adaptive application of the local Hadamard rotation, $H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$. This enables the definition of two subroutines, P1 and P2, which operate with and without the Hadamard gate, respectively, each targeting the detection of different types of errors. These subroutines can be iteratively alternated or optimally selected to further enhance the performance of the protocols.
@@
that accounts for the resource overheads required to complete the protocol, i.e., it evaluates the efficiency of the protocol. Rigorously, the value of the amount of copies obtained is restricted to an accuracy condition, so that the protocol is considered successful if the final fidelity satisfies $F>1-\epsilon$. For recurrence protocols with a single copy as output, the obtained number of copies is given by $\frac{\prod_{k}P_{k}}{m^{k}}$, where $P_{k}$ is the success probability of the $k^{\rm th}$ iteration and $m$ is the number of copies used in each round.
@@
2.4 Superposition of quantum channels
@@
Superposition of different tasks is achieved by introducing a quantum control register in a superposition state, which acts as a control to exchange roles of states or tasks via a controlled-SWAP operation, also known as Fredkin gate [59, 60], which is a non-Clifford operation described as
@@
3 Superposed EPP: role exchange
@@ 3.1 Basic protocol
    $\left|\Phi_{00}\right\rangle\left\langle\Phi_{00}\right|_{\rm A_{\rm c}B_{\rm c}}$ (13)
    $\otimes\left(\sum_{i,j=0}^{1}{p}_{i,j}\Sigma_{i,j}\left|\Phi_{00}\right\rangle\left\langle\Phi_{00}\right|_{\rm A_{1}B_{1}}\Sigma^{\dagger}_{i,j}\right)$
@@
    where we can assume without loss of generality that $\Sigma_{ij}\equiv\mathbb{1}_{\rm A}\otimes{\sigma_{ij}}_{\rm B}$, where $\{\sigma_{00},\sigma_{01},\sigma_{10},\sigma_{11}\}=\{\mathbb{1},\sigma_{z},\sigma_{x},\sigma_{y}\}$, such that $\Sigma_{ij}\left|\Phi_{00}\right\rangle=\left|\Phi_{ij}\right\rangle$ maps all the basis Bell states. Here ${p}_{i,j}$ determines the strength of each noisy contribution, where ${p}_{00}$ corresponds to the state initial fidelity.
@@
    $\left|{ss}\right\rangle\left\langle{ss}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{i,j}\right\rangle\left\langle\Phi_{i,j}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{q,r}\right\rangle\left\langle\Phi_{q,r}\right|_{\rm A_{2}B_{2}}\right)$ (14)
@@
    $+$ $\frac{1}{2}\left|{11}\right\rangle\left\langle{00}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{q,r}\right\rangle\left\langle\Phi_{i,j}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{i,j}\right\rangle\left\langle\Phi_{q,r}\right|_{\rm A_{2}B_{2}}\right),$
    where $\left|{ss}\right\rangle\left\langle{ss}\right|$ refers to both the $\left|{00}\right\rangle\left\langle{00}\right|$ and $\left|{11}\right\rangle\left\langle{11}\right|$ terms.
@@
    $\left|{ss}\right\rangle\left\langle{ss}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{i\oplus q,j}\right\rangle\left\langle\Phi_{i\oplus q,j}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{q,r\oplus j}\right\rangle\left\langle\Phi_{q,r\oplus j}\right|_{\rm A_{2}B_{2}}\right)$ (15)
    $+$ $\frac{1}{2}\left|{00}\right\rangle\left\langle{11}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{i\oplus q,j}\right\rangle\left\langle\Phi_{q\oplus i,r}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{q,r\oplus j}\right\rangle\left\langle\Phi_{i,j\oplus r}\right|_{\rm A_{2}B_{2}}\right)$
    $+$ $\frac{1}{2}\left|{11}\right\rangle\left\langle{00}\right|_{\rm A_{\rm c}B_{\rm c}}\otimes\left(\sum_{i,j,q,r=0}^{1}{p}_{i,j}\left|\Phi_{q\oplus i,r}\right\rangle\left\langle\Phi_{i\oplus q,j}\right|_{\rm A_{1}B_{1}}\otimes{p}_{q,r}\left|\Phi_{i,j\oplus r}\right\rangle\left\langle\Phi_{q,r\oplus j}\right|_{\rm A_{2}B_{2}}\right).$
@@
    Subsequently, another controlled-SWAP (cSWAP) operation can be implemented, although this step is not necessary. Finally, the second copy is measured locally in the $Z$ basis at each site. If parties $A$ and $B$ obtain the same outcome ($j\oplus r=0$), the first state is retained. The only components that survive the measurement are then the ones coming from $\left({p}_{q,j}\left|\Phi_{i,0}\right\rangle\left\langle\Phi_{q,0}\right|_{\rm A_{2}B_{2}}\right)$. Simultaneously, the control state is measured locally in the $X$ basis by $\rm A$ and $\rm B$, post-selecting the desired outcomes as stated in the following.
@@
where $P=\sum_{j}{p}^{2}_{jj}$ is a normalization constant that equals the success probability, and where the new fidelity is $F=\dfrac{\sum_{i}{p}^{2}_{i0}}{\sum_{j}{p}^{2}_{jj}}$.
@@
Input: Three copies of a noisy bipartite entangled state $\rho_{0}$. Alice and Bob perform identical operations.
  1. 1.
    Prepare three copies of a noisy entangled state $\rho_{0}$.
  2. 2.
    Apply suitable bilateral rotations $(H,U_{0})$ depending on the protocol variant (Oxford, P1, P2). This can be optimized such that the best variant is chosen in each iteration. Then, bilaterally apply a cSWAP operation with the first copy acting as superposition-control.
  3. 3.
    Apply a bilateral cNOT operation between the second and third copy.
  4. 4.
    After a second cSWAP for recombination, measure the third copy in the Z basis and the control copy in the X basis. Keep the second copy if outcomes coincide for both measurements.
Output: A purified entangled state with enhanced fidelity. A basic EPP protocol is executed but with the control and target roles coherently exchanged during the process.
Algorithm 1 Superposed EPP. Role exchange
To address this, we turn our attention to a realistic scenario: a $3\rightarrow 1$ protocol involving a control copy identically noisy as the others, as analyzed in Fig. 3. We show how the role-exchanging superposed entanglement purification approach introduced above, can significantly surpass single selection protocols [17, 23, 18], and also more elaborated ones such as the double selection protocol, Fig. 1, [41], which is considered among the most efficient EPPs modifications using $3$ raw EPR copies [43]. We consider different initial entangled states, affected by amplitude damping in Fig. 3 (a), and dephasing noise in Fig. 3 (b). In addition, we consider the P1-or-P2 variants of all the protocols (see Figs. 1 - 2), meaning that the best subroutine (P1 or P2) is selected in each iteration. We analyze the performance for different kinds of initial states. The plots show the increase in fidelity $F$ (or decrease in infidelity, defined as $1-F$) with the number of steps, where one observes that the superposed protocol performs better than the elementary ones.
We provide in Appendix Appendix A. Density matrix evolution for different EPP protocols a detailed analysis of the density matrix evolution of the superposed EPP protocol compared with the Oxford and double selection strategies, for different instances of initial states, including single Pauli noise and Werner states.
@@
Fig. 3: (a) Fidelity evolution as a function of the number iterations for the BBPSSW P1-or-P2 EPPs with initial states affected (both sides) by amplitude damping noise of strength $q=0.6$. (b) Infidelity evolution as a function of the number iterations for the different P1-or-P2 Oxford variation protocols with initial states affected by dephasing noise and fidelity $F_{0}=0.6$ (see Appendix Appendix A. Density matrix evolution for different EPP protocols for a detailed analysis of each protocol for correcting Pauli noises). In all cases, depolarizing the initial states into Werner form decreases the overall protocols performances.
3.2 Imperfect operations